# The Bandit Whisperer: Communication Learning for Restless Bandits

### v1

@@ Abstract
Applying Reinforcement Learning (RL) to Restless Multi-Arm Bandits (RMABs) offers a promising avenue for addressing allocation problems with resource constraints and temporal dynamics. However, classic RMAB models largely overlook the challenges of (systematic) data errors—a common occurrence in real-world scenarios due to factors like varying data collection protocols and intentional noise for differential privacy. We demonstrate that conventional RL algorithms used to train RMABs can struggle to perform well in such settings. To solve this problem, we propose the first communication learning approach in RMABs, where we study which arms, when involved in communication, are most effective in mitigating the influence of such systematic data errors. In our setup, the arms receive Q-function parameters from similar arms as messages to guide behavioral policies, steering Q-function updates. We learn communication strategies by considering the joint utility of messages across all pairs of arms and using a Q-network architecture that decomposes the joint utility. Both theoretical and empirical evidence validate the effectiveness of our method in significantly improving RMAB performance across diverse problems.
@@ 1 Introduction
Related works. To the best of our knowledge, this is the first work that introduces communication into RMABs. This is perhaps surprising given that communication is an important topic that has been extensively studied in other multi-agent problems. In Dec-POMDPs (Oliehoek, Amato et al. 2016; Kwak et al. 2010; Nair et al. 2004), previous works explore decentralized communication learning (Sukhbaatar, Fergus et al. 2016; Hoshen 2017; Jiang and Lu 2018; Singh, Jain, and Sukhbaatar 2019; Wang et al. 2019b), the emergence of natural language (Foerster et al. 2016; Das et al. 2017; Lazaridou, Peysakhovich, and Baroni 2017; Mordatch and Abbeel 2018; Kang, Wang, and de Melo 2020), and implicit communication by coordination graphs (Böhmer, Kurin, and Whiteson 2020; Wang et al. 2021; Kang et al. 2022; Yang et al. 2022), influence on each other (Jaques et al. 2019; Wang et al. 2019a), and attention mechanisms (Huang, Mordatch, and Pathak 2020; Kurin et al. 2020; Dong et al. 2022, 2023). Different from these works, we consider communication under noisy returns and in the RMAB setting where the central planner policy constrained by resource budgets determines the influence of communication.
@@ 2.1 RMABs with Noisy Arms
$\mathcal{A}_{i}$ is a finite set of discrete actions. Each action $a\in\mathcal{A}_{i}$ has a cost $\mathcal{C}_{i}(a)$. Following the standard bandit assumption that an arm can be ”not pulled” at no cost, we set $\mathcal{C}_{i}(0)=0$. The probability of transitioning from one state to another given an action is specified by the function $T_{i}:\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\rightarrow[0,1]$. In line with previous work (Zhao et al. 2023), we assume that $\mathcal{S}_{i},\mathcal{A}_{i}$, and $\mathcal{C}_{i}$ are the same for all arms $i\in[N]$ and omit the subscript $i$ for simplicity. Each arm has a feature vector $\bm{z}_{i}\in\mathbb{R}^{m}$ that provides useful information about the transition dynamics of arm $i$. We define the arm with the most similar features with arm $i$ as $\nu_{i}=\arg\min_{j}\|z_{j}-z_{i}\|_{2}$.
@@ 2.2 Reinforcement Learning in RMABs
Deep reinforcement learning has been proven to be efficient in solving the RMAB problem (Avrachenkov and Borkar 2022; Killian et al. 2022; Zhao et al. 2024). To learn arm policies, we take the Lagrangian relaxation of the budget constraint (Hawkins 2003), defining
@@ 4 Theoretical Analysis
as the lowest probability of any state-action pair under the stationary distribution $\mu_{\pi_{\nu_{i}}}$ induced by the behavior policy $\pi_{\nu_{i}}$ on the receiver MDP, where $\nu_{i}$ is the arm with the most similar feature to arm $i$. We further define the mixing time (Hsu, Kontorovich, and Szepesvári 2015; Li et al. 2020),
$t_{\texttt{mix}}=\min\left\{t|\max_{s_{0},a_{0}}\sup_{s,a}|P^{t}(s,a|s_{0},a_{0})-\mu_{\pi_{\nu_{i}}}(s,a)|\leq\frac{1}{4}\right\},$ (11)
@@
The Proof is detailed in Appendix C.2. Proposition 2 shows that communication can accelerate arm $i$’s Q-learning as long as $\pi_{\nu_{i}}$ visits every state-action pair with a small probability and is not too slowly disentangling from the initial state. This finding is solely dependent upon the characteristics of the behavior policy in the receiver MDP $\mathcal{M}_{i}$ and does not require the sender or the receiver to be non-noisy or noisy.
@@ 5 Experiments
SIS Epidemic Model (Yaesoubi and Cohen 2011; Killian et al. 2022; Zhao et al. 2024): Each arm $p$ represents a subpopulation in a geographic region, with the number of uninfected individuals being the state $s_{p}$. We implement a budget-constrained intervention framework with three actions: $a_{0}$ represents no intervention, $a_{1}$ requires weekly COVID testing, and $a_{2}$ involves distributing face masks. Actions change the disease transmission parameters in a targeted manner, reflecting their real-world public-health impacts. The reward noise is due to variability in data collection, which may be caused by poor COVID testing equipment (Embrett et al. 2022; Schmitt-Grohé, Teoh, and Uribe 2020) and influence of authorities or impacted populations (keeping the COVID numbers low) (Paulus et al. 2023).
@@
Table 1: Performance (return averaged across 100 random seeds) with increased noise levels, normalized with respect to $\mathtt{No\ Noise}$ and $\mathtt{Comm}$ (ours) before communication starts. We test Gaussian Noise $\mathcal{N}(0,\sigma)$, and the $x\%$ noise level means $\sigma=x/100$.
@@
Table 2: Performance (return averaged across 100 random seeds) with numbers of noisy arms, normalized with respect to $\mathtt{No\ Noise}$ and $\mathtt{Comm}$ (ours) before communication starts.
@@
Robustness.  Tables 1 and  2 show the performance of different methods under various settings (see Appendix for more ablations). In $\mathtt{No\ Noise}$, all arms receive accurate, true reward signals without any noise. Arms learn individually on their own data without any communication. In $\mathtt{Random\ Comm}$, each arm receives communication from a randomly chosen arm. The results demonstrate that our method is robust under increasing noise levels, numbers of noisy arms, and different noise distributions. In contrast, the performance of baselines $\mathtt{No\ Comm}$ and $\mathtt{Fixed\ Comm}$ changes significantly when the environment changes. In particular, on Synthetic, when the number of noisy arms increases from 15 to 17 (Table 2), our $\mathtt{Comm}$ method has only 3% drop in performance, while $\mathtt{Fixed\ Comm}$ has 42 percentage points drop in performance, despite that $\mathtt{Fixed\ Comm}$ uses ground truth information about which arms are noisy that is not known to our $\mathtt{Comm}$ method. Similar trends can be seen on SIS and ARMMAN. These results demonstrate that, compared to a fixed communication protocol with access to oracle information, our approach for communication learning provides better robustness. Additionally, the performance of $\mathtt{No\ Comm}$ supports the argument that conventional RL methods struggle in noisy RMABs.
@@ 6 Closing Remarks
The study of RMABs in environments with noisy and unreliable data sources underscores the need for innovative solutions in RL methods. We introduce a novel communication learning algorithm to alleviate the influence of noisy data so that the central planner can make better coordination decisions. Theoretically, we prove that communication can help reduce errors in value function approximation if the behavioral policy’s state-action occupancy probability is lower bounded by a suitable constant. Empirically, we show that our approach significantly improves RMAB performance. Our findings advance the understanding of RMAB dynamics and enhance their applicability across real-world settings.
@@ Appendix B Additional Experiment Details and Results
Table 3: Performance (return averaged across 100 random seeds) with various noise distributions, normalized with respect to $\mathtt{No\ Noise}$ and $\mathtt{Comm}$ (ours) before communication starts. Gaussian noise is $\mathcal{N}(0,1)$, Uniform noise is $U(-0.5,0.5)$, and Mixture noise has 50% chance of being the Gaussian and 50% chance of being the Uniform. We abbreviate $\mathtt{Nearest\ Neighbor\ Comm}$ as $\mathtt{NN\ Comm}$
@@ Appendix D Additional Related Works
RMAB have been widely used to model multiple interacting agents (Jaques et al. 2019; Yu et al. 2022; Shapley 1953; Littman 1994; Jin et al. 2021; Xie et al. 2020; Behari et al. 2024). For binary action RMABs, Whittle index based Q-learning approaches have been investigated (Nakhleh et al. 2021; Xiong and Li 2023; Fu et al. 2019; Avrachenkov and Borkar 2022). Recent works also studied deep RL approaches for multi-action RMABs (Killian et al. 2021; Zhao et al. 2024; Killian et al. 2022). However, existing works on multi-action RMABs largely overlook the challenge of data errors, which is common in real-world scenarios. Our novel communication learning approach enables arms to mitigate the influence of systematic data errors.

### v3

@@ Abstract
Applying Reinforcement Learning (RL) to Restless Multi-Arm Bandits (RMABs) offers a promising avenue for addressing allocation problems with resource constraints and temporal dynamics. However, classic RMAB models largely overlook the challenges of (systematic) data errors, a common occurrence in real-world scenarios due to factors like varying data collection protocols and intentional noise for differential privacy. We demonstrate that conventional RL algorithms used to train RMABs can struggle to perform well in such settings. To solve this problem, we propose the first communication learning approach in RMABs, where we study which arms, when involved in communication, are most effective in mitigating the influence of such systematic data errors. In our setup, the arms receive Q-function parameters from similar arms as messages to guide behavioral policies, steering Q-function updates. We learn communication strategies by considering the joint utility of messages across all pairs of arms and using a Q-network architecture that decomposes the joint utility. Both theoretical and empirical evidence validate the effectiveness of our method in significantly improving RMAB performance across diverse problems.
@@ 1 Introduction
Related works. To the best of our knowledge, this is the first work that introduces communication into RMABs. This is perhaps surprising given that communication is an important topic that has been extensively studied in other multi-agent problems. In Dec-POMDPs (Oliehoek, Amato et al. 2016; Kwak et al. 2010; Nair et al. 2004), previous works explore decentralized communication learning (Sukhbaatar, Fergus et al. 2016; Hoshen 2017; Jiang and Lu 2018; Singh, Jain, and Sukhbaatar 2019; Wang et al. 2019b), the emergence of natural language (Foerster et al. 2016; Das et al. 2017; Lazaridou, Peysakhovich, and Baroni 2017; Mordatch and Abbeel 2018; Zhao et al. 2024b; Kang, Wang, and de Melo 2020), and implicit communication by coordination graphs (Böhmer, Kurin, and Whiteson 2020; Wang et al. 2021; Kang et al. 2022; Yang et al. 2022), influence on each other (Jaques et al. 2019; Wang et al. 2019a), and attention mechanisms (Huang, Mordatch, and Pathak 2020; Kurin et al. 2020; Dong et al. 2022, 2023). Different from these works, we consider communication under noisy returns and in the RMAB setting where the central planner policy constrained by resource budgets determines the influence of communication.
@@ 2.1 RMABs with Noisy Arms
$\mathcal{A}_{i}$ is a finite set of discrete actions. Each action $a\in\mathcal{A}_{i}$ has a cost $\mathcal{C}_{i}(a)$. Following the standard bandit assumption that an arm can be ”not pulled” at no cost, we set $\mathcal{C}_{i}(0)=0$. The probability of transitioning from one state to another given an action is specified by the function $T_{i}:\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\rightarrow[0,1]$. In line with previous work (Zhao et al. 2023; Elmachtoub, Gupta, and Zhao 2023; Elmachtoub et al. 2023; Zhao et al. 2022) we assume that $\mathcal{S}_{i},\mathcal{A}_{i}$, and $\mathcal{C}_{i}$ are the same for all arms $i\in[N]$ and omit the subscript $i$ for simplicity. Each arm has a feature vector $\bm{z}_{i}\in\mathbb{R}^{m}$ that provides useful information about the transition dynamics of arm $i$. We define the arm with the most similar features with arm $i$ as $\nu_{i}=\arg\min_{j}\|z_{j}-z_{i}\|_{2}$.
@@ 2.2 Reinforcement Learning in RMABs
Deep reinforcement learning has been proven to be efficient in solving the RMAB problem (Avrachenkov and Borkar 2022; Killian et al. 2022; Zhao et al. 2024a). To learn arm policies, we take the Lagrangian relaxation of the budget constraint (Hawkins 2003), defining
@@ 4 Theoretical Analysis
as the lowest probability of any state-action pair under the stationary distribution $\mu_{\pi_{\nu_{i}}}$ induced by the behavior policy $\pi_{\nu_{i}}$ on the receiver MDP, where $\nu_{i}$ is the arm with the most similar feature to arm $i$. We further define the mixing time (Hsu, Kontorovich, and Szepesvári 2015; Li et al. 2020) $t_{\texttt{mix}}:=$
$\min\left\{t|\max_{s_{0},a_{0}}\sup_{s,a}|P^{t}(s,a|s_{0},a_{0})-\mu_{\pi_{\nu_{i}}}(s,a)|\leq\frac{1}{4}\right\},$ (11)
@@
The Proof is detailed in Appendix C.2 Proposition 2 shows that communication can accelerate arm $i$’s Q-learning as long as $\pi_{\nu_{i}}$ visits every state-action pair with a small probability and is not too slowly disentangling from the initial state. This finding is solely dependent upon the characteristics of the behavior policy in the receiver MDP $\mathcal{M}_{i}$ and does not require the sender or the receiver to be non-noisy or noisy.
@@ 5 Experiments
SIS Epidemic Model (Yaesoubi and Cohen 2011; Killian et al. 2022; Zhao et al. 2024a): Each arm $p$ represents a subpopulation in a geographic region, with the number of uninfected individuals being the state $s_{p}$. We implement a budget-constrained intervention framework with three actions: $a_{0}$ represents no intervention, $a_{1}$ requires weekly COVID testing, and $a_{2}$ involves distributing face masks. Actions change the disease transmission parameters in a targeted manner, reflecting their real-world public-health impacts. The reward noise is due to variability in data collection, which may be caused by poor COVID testing equipment (Embrett et al. 2022; Schmitt-Grohé, Teoh, and Uribe 2020) and influence of authorities or impacted populations (keeping the COVID numbers low) (Paulus et al. 2023).
@@
Table 1: Average returns (100 seeds) with increased noise, normalized with respect to $\mathtt{No\ Noise}$ and $\mathtt{Comm}$ (ours) before communication (see Remark 7 in Appendix B). We test Gaussian Noise $\mathcal{N}(0,\sigma)$. $x\%$ noise level means $\sigma=x/100$.
@@
Table 2: Performance (return averaged across 100 random seeds) with numbers of noisy arms, normalized with respect to $\mathtt{No\ Noise}$ and $\mathtt{Comm}$ (ours) before communication starts (see Remark 7 in Appendix B).
@@
Robustness.  Tables 1 and  2 show the performance of different methods under various settings (see Appendix B for more ablations). In $\mathtt{No\ Noise}$, all arms receive accurate, true reward signals without any noise. Arms learn individually on their own data without any communication. In $\mathtt{Random\ Comm}$, each arm receives communication from a randomly chosen arm. The results demonstrate that our method is robust under increasing noise levels, numbers of noisy arms, and different noise distributions. In contrast, the performance of baselines $\mathtt{No\ Comm}$ and $\mathtt{Fixed\ Comm}$ changes significantly when the environment changes. In particular, on Synthetic, when the number of noisy arms increases from 15 to 17 (Table 2), our $\mathtt{Comm}$ method has only 3% drop in performance, while $\mathtt{Fixed\ Comm}$ has 42 percentage points drop in performance, despite that $\mathtt{Fixed\ Comm}$ uses ground truth information about which arms are noisy that is not known to our $\mathtt{Comm}$ method. Similar trends can be seen on SIS and ARMMAN. These results demonstrate that, compared to a fixed communication protocol with access to oracle information, our approach for communication learning provides better robustness. Additionally, the performance of $\mathtt{No\ Comm}$ supports the argument that conventional RL methods struggle in noisy RMABs.
@@ 6 Closing Remarks
The study of RMABs in environments with noisy and unreliable data sources underscores the need for innovative solutions in RL methods. We introduce a novel communication learning algorithm to alleviate the influence of noisy data so that the central planner can make better coordination decisions. Theoretically, we prove that communication can help reduce errors in value function approximation if the behavioral policy’s state-action occupancy probability is lower bounded by a suitable constant. Empirically, we show that our approach significantly improves RMAB performance.
Acknowledgments
The work is supported by Harvard HDSI funding
@@ Appendix B Additional Experiment Details and Results
Table 3: Performance (return averaged across 100 random seeds) with various noise distributions, normalized with respect to $\mathtt{No\ Noise}$ and $\mathtt{Comm}$ (ours) before communication starts. Gaussian noise is $\mathcal{N}(0,1)$, Uniform noise is $U(-0.5,0.5)$, and Mixture noise has 50% chance of being the Gaussian and 50% chance of being the Uniform. We abbreviate $\mathtt{Nearest\ Neighbor\ Comm}$ as $\mathtt{NN\ Comm}$.
@@
Our method is robust to both biased and unbiased reward errors. In Synthetic, with biased noise $\mathcal{N}(z,1)$ across varying $z$, our method consistently outperforms baselines, as shown in Table 5.
Synthetic (N=21, B=15)
Method $\mathcal{N}(0,1)$ $\mathcal{N}(0.1,1)$ $\mathcal{N}(0.2,1)$
$\mathtt{Comm}$ 75% 64% 56%
$\mathtt{No\ Comm}$ 62% 50% 43%
$\mathtt{Fixed/Comm}$ 21% 29% 29%
$\mathtt{NN\ Comm}$ 44% 33% 24%
Table 5: Performance (return averaged across 100 random seeds) with biased noise $\mathcal{N}(z,1)$ across varying $z$, normalized with respect to $\mathtt{No\ Noise}$ and $\mathtt{Comm}$ (ours) before communication starts. We abbreviate $\mathtt{Nearest\ Neighbor\ Comm}$ as $\mathtt{NN\ Comm}$
Remark 7.
Notice that in Tables 1, 2, 3, and 5, we normalize the returns of all algorithms under consideration by
$\frac{\text{return}-\text{return($\mathtt{Comm}$ at epoch 200)}}{\text{return($\mathtt{No\ Noise}$ at epoch 200})-\text{return($\mathtt{Comm}$ at epoch 200})}.$
Communication learning starts at epoch 200.
@@ Appendix D Additional Related Works
RMAB and RL in general have been widely used to model multiple interacting agents (Jaques et al. 2019; Yu et al. 2022; Shapley 1953; Littman 1994; Jin et al. 2021; Xie et al. 2020; Behari et al. 2024; Zhao 2023; Choromanski et al. 2023; Seow et al. 2024; Sehanobish et al. 2023; Boehmer et al. 2024) For binary action RMABs, Whittle index based Q-learning approaches have been investigated (Nakhleh et al. 2021; Xiong and Li 2023; Fu et al. 2019; Avrachenkov and Borkar 2022). Recent works also studied deep RL approaches for multi-action RMABs (Killian et al. 2021; Zhao et al. 2024a; Killian et al. 2022). However, existing works on multi-action RMABs largely overlook the challenge of data errors, which is common in real-world scenarios. Our novel communication learning approach enables arms to mitigate the influence of systematic data errors.